• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name danez

    Log entry time 16:00:27 on February 08, 2011

    Entry number 346161

    keyword=Left HRS Electronics

    Like the right arm, the left arm's ADC gates were corrected (with Vince) and the 50ns gate for BPM was created and sent to slot 25 of ROC4.

    The L1A signal did not need to be split, since it was already split when it goes through its delay. (The right arm doesn't have a delay yet, since the timing hasn't been done yet, since the RG8s aren't connected yet.)

    The scaler map was updated for the left arm, including a copy of the 100kHz pulser.